### v1

# Karma: An Experimental Study

Abstract.
A system of non-tradable credits that flow between individuals like karma, hence proposed under that name, is a mechanism for repeated resource allocation that comes with attractive efficiency and fairness properties, in theory. In this study, we test karma in an online experiment in which human subjects repeatedly compete for a resource with time-varying and stochastic individual preferences or urgency to acquire the resource. We confirm that karma has significant and sustained welfare benefits even in a population with no prior training. We identify mechanism usage in contexts with sporadic high urgency, more so than with frequent moderate urgency, and implemented as an easy (binary) karma bidding scheme as particularly effective for welfare improvements: relatively larger aggregate efficiency gains are realized that are (almost) Pareto superior. These findings provide guidance for further testing and for future implementation plans of such mechanisms in the real world.
conference: ; ;
1Automatic Control Lab, ETH Zurich. Emails: {elokdae,bsaverio,dorfler}@ethz.ch
2Department of Sociology, University of Zurich. Email: heinrich.nax@uzh.ch
Research supported by NCCR Automation, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number $180545$).
1. Introduction
Efficiency and fairness in determining who gets what and when are the two major objectives of economists thinking about scarcity, and many interesting matchmaking and market design solutions have been proposed (Roth, 2015). An important class of allocation problems is when goods are repeatedly allocated in the population. In situations like these, it is often the case that one person relative to another gets higher utility from the good in one period but lower utility in another (which we shall capture by the notion of urgency). A recent innovation to address exactly those kinds of allocation problems goes under the name karma (Vishnumurthy et al., 2003; Elokda et al., 2023; Vuppalapati et al., 2023). The karma mechanism mirrors what Western popular culture associates with the notions of karma and samsara stemming from Indian religions according to which one’s deeds in the present (karma) affect the quality of one’s future life (phala) and thereafter (samsara) (Reichenbach, 1990; Kyabgon, 2015). The mechanism is implemented via individual accounts of non-tradable credits called karma. Individuals may bid some amount from their current account at every instance of resource allocation. The highest bidders win, obtain priority for the resource, pay their bids, which are then transferred to the accounts of those who did not win.
Such mechanisms have been proposed previously in non-human contexts to manage scarcity in, for example, traffic (Elokda et al., 2022), transportation (Censi et al., 2019; Salazar et al., 2021), cloud computing (Vuppalapati et al., 2023), and peer-to-peer networks (Vishnumurthy et al., 2003). In these works, agents are assumed to follow perfect Nash equilibrium as they can be programmed to do so. In human contexts, the most well-known example of a karma-like mechanism is the ‘choice system’ adopted by U.S. food banks (Prendergast, 2022), where the role of karma tokens is taken by ‘shares.’ As with karma, food banks under the choice system bid shares to obtain priority for available food donations, and the total payment of shares is redistributed at the end of each day. This real-world implementation of a karma-like system indeed resulted in significant aggregate gains in terms of food bank participation and fluidity of donations. In this paper, we move to a controlled test of such a mechanism as the behavioral uptake by the human actors involved as well as the distributional consequences continue to remain unstudied experimentally.
Karma and the aforementioned choice system are different from other token-based mechanisms that have been implemented and proposed in the past, including those that use tradable credits (Cramton and Kerr, 2002; Sovacool, 2011; Yang and Wang, 2011), non-tradable vouchers (Daniels and Trebilcock, 2013; Hsieh and Urquiola, 2006; Kim et al., 2021), or so-called artificial currencies or scrips (Johnson et al., 2014; Gorokh et al., 2021a, b; Banerjee et al., 2023). In these other mechanisms, some type of voucher or token is typically issued periodically, e.g., at the beginning of every week or month, to be consumed over the period. Karma offers three distinguishing features in comparison to these mechanisms. First, karma is non-tradable even in a black market, as it is associated to individual accounts and can be transferred only in fixed resource allocation instances. Second, one karma token does not equal one resource unit; the price instead arises dynamically from bidding. This has the dual benefit of making the system indifferent to the exact amount of karma in it, and enabling individuals to express their time-varying private urgency. Third, karma forms a closed economy: it is transferred among individuals in a manner that preserves the karma per capita over time. This is advantageous for both system and individuals. From the system perspective, the karma mechanism can be infinitely repeated with minimal central intervention, as no central endowment of karma is needed other than the initial endowment. From the individual perspective, karma is forgiving as it enables individuals to immediately replenish their accounts through yielding, rather than waiting a potentially very long time for the next central endowment.
Figure 1 illustrates how karma works out to the benefit of everyone at hand of an example involving three intersection encounters. Let’s start with the encounter in the top center. The high-urgency lila car has a current karma account of 9 and bids 4, thus outbidding the low-urgency blue car whose karma account is also 9 but bids 2. As a result, the blue car’s karma account goes up by 4 to 13. Let’s move along clockwise. Blue now happens to have high urgency and bids 4, thus outbidding and getting priority over orange whose karma account is nine and bids 3. Now the orange karma goes up by 4 to 13. Let’s move along clockwise. Orange now has high urgency and bids 4, thus outbidding low-urgency lila who has 5 karma left and bids 1. Thus the circle closes, et cetera, et cetera.
Figure 1. Dynamic resource allocation with karma.
The example illustrates that karma is similar in spirit to the ‘universal currency’ of reputation (Milinski, 2016), but makes the direct linkages and transfers between users’ karma accounts explicit, while reputation is something that tends to exist in humans’ individual perceptions of one another informally. Similar to how ‘image scoring,’ which is an example of an explicit reputation mechanism, formalizes the notion of indirect reciprocity in reputation-based interactions (Nowak and Sigmund, 1998), karma can be viewed as the formal equivalent of ‘fair turn-taking’ in repeated resource allocation tasks. While image scoring has been subjected to the behavioral test in the past (Wedekind and Milinski, 2000), it remains to do so for karma.
Further, the example also illustrates another feature of karma, as was mentioned above, that distinguishes karma from fixed-value tokens and from other forms of rules ensuring turn-taking: it allows individuals to dynamically express how intense their private preference for the resource is. Clearly, such a mechanism may be advantageous when preference intensities/urgencies in the population are heterogeneous and not perfectly correlated over time. Indeed, (Elokda et al., 2023) show that farsighted Nash equilibria are almost fully efficient with respect to the private urgency of the users.
To date, the literature on karma focused on the theoretical properties of the mechanism compared with random allocation and other domain-specific schemes such as dynamic pricing or max-min allocation (Vuppalapati et al., 2023; Elokda et al., 2022, 2023), all predicted on idealized assumptions such as rationality, far-sightedness and perfect adoption. In this paper, we complement the pre-existent literature by a first behavioral investigation with real humans who do not necessarily follow such idealized assumptions. We want to know whether humans find it natural to adapt the mechanism, and whether inexperienced players are able to realize efficiency gains. To this end, we conducted a controlled experiment to better understand whether and how effectively karma might be implemented in human populations as a resource governance mechanism. In our analysis we focus on the overall efficiency effects of karma as well as on distributional consequences and associated fairness properties of the mechanism. In our treatments, we vary the nature of the dynamic preference intensity distributions (more frequent and less intense versus less frequent and more intense) and the auction process of the karma mechanism (binary bidding versus full bidding). We use random allocation as a benchmark to compare our findings to. This domain-independent benchmark is representative of common schemes that are unaware of the individual private preferences, including fixed turn-taking schemes and schemes using fixed-value tokens.
Our main findings summarize as follows. Overall, compared with random allocation, karma leads to substantial efficiency gains, and these gains benefit almost all participants. In fact, only non-adopters, that is, individuals who do not participate in karma bidding actively themselves, do not benefit and may be worse off. The treatment variations suggest that benefits are particularly pronounced in situations when preference intensities are dynamically more intense and less frequent, and the bidding scheme is designed to be minimal (i.e., binary). These findings provide a first benchmark that karma may be used beneficially in human interactions. Our study also points in several directions for further theoretical and experimental investigation.
2. Experimental Methods
We conducted a balanced two-by-two factorial experiment with 400 subjects in total. Treatments varied in the dynamic urgency process of the participants and the richness of the karma scheme. For urgency, we distinguished between a low stake process where participants have frequent events with moderate urgency, and a high stake process where participants have rare events with high urgency. For richness, we tested a binary scheme where participants can choose from only two bid levels that depend on their karma, and a full range scheme where participants have full choice over the bid up to their karma.
2.1. The Game
The game we study is one that proceeds with $N$ participants over $T$-many rounds. All participants receive an initial endowment of karma $k(0)=k(1)={k_{\textup{init}}}\in\mathbb{N}$ and an initial game score $s(0)=0$. At each round $t\in\{1,\dots,T\}$, participants are randomly matched in pairs to compete over a shared resource. Each participant is given an urgency value $u(t)\in\{{u_{\textup{l}}},\>{u_{\textup{h}}}\}$ that is drawn randomly and independently from a process $\mathbb{P}(u)$ that is identical for all players, and must place a bid $b(t)\leq k(t)$. In each pairwise matching, the higher bidder gets allocated the resource: the score $s(t)=s(t-1)+u(t)$ increases by the urgency and the bid is collected as payment. The lower bidder does not get allocated the resource: the score $s(t)=s(t-1)$ does not increase and no payment is collected. Ties are settled randomly. To keep the total amount of karma constant, the total payment ${p_{\textup{tot}}}$ is uniformly redistributed to the participants in an integer preserving manner. In case ${p_{\textup{tot}}}$ is not divisible by $N$, $\left\lfloor\frac{{p_{\textup{tot}}}}{N}\right\rfloor$ is redistributed to everyone, and a random subset of ${p_{\textup{tot}}}-N\left\lfloor\frac{{p_{\textup{tot}}}}{N}\right\rfloor$ participants receive one additional karma. Moreover, each participant is allowed a maximum karma level ${k_{\textup{max}}}$ that the redistribution respects. Participants with ${k_{\textup{max}}}$ do not receive additional redistribution which instead gets issued uniformly to the others. After the redistribution the game proceeds to the next round $t+1$.
Table 3 in Appendix A summarizes the notation introduced above.
2.2. Treatments
We follow a 2x2 factorial treatment design, where we vary the dynamic urgency process of the participants and the richness of the karma scheme. Table 1 summarizes the treatment configurations schematically. For each of the four resulting treatments we run 5 independent game experiments with 20 participants per experimental game resulting in 100 subjects per treatment and 400 subjects in total.
Table 1. 2x2 experimental design.
Richness of scheme
Binary Full range
Urgency process
Low stake [Uncaptioned image]    [Uncaptioned image] [Uncaptioned image]    [Uncaptioned image]
High stake [Uncaptioned image]    [Uncaptioned image] [Uncaptioned image]    [Uncaptioned image]
2.2.1. Urgency Process
The urgency processes in both treatment variations of low stake and high stake have the same low urgency ${u_{\textup{l}}}=1$, but differ in the magnitude and frequency of the high urgency (low stake: ${u_{\textup{h}}}=5$, ${\mathbb{P}}({u_{\textup{h}}})=0.5$; high stake: ${u_{\textup{h}}}=9$, ${\mathbb{P}}({u_{\textup{h}}})=0.25$). Both processes have the same urgency on average $\mathbb{E}(u)=3$, and therefore the same expected scores under random allocation. Low stake represents the case where the high urgency event is frequent but moderate, while in high stake the high urgency event is rare but severe.
2.2.2. Richness of Karma Scheme
In the treatment variation of binary, participants can only choose between two bid levels: 0 or $\left\lfloor\frac{k}{2}\right\rfloor$ (i.e., half their karma, rounded down to the nearest integer). In the treatment variation of full range, participants can choose any integer bid up to their karma. The binary treatment aims at investigating the behavioral effects of a reduced action space, whereby subjects either bid or not. Using tools from (Elokda et al., 2023), we designed the urgency processes in such a way that there do not result noteworthy efficiency differences associated with the Nash equilibria of the two schemes. Therefore, any differences observed in terms of outcomes between the two bidding schemes will be due to behavioral effects.
2.3. Experimental Implementation
The game is implemented as a real-time online experiment using oTree (Chen et al., 2016). Participants are given the opportunity to familiarize with the game over $5$ rounds that do no contribute to the game score, after which the main game proceeds with $T=50$ rounds. In each round, participants get drawn a random urgency. On the decision page, they must place a bid, and then a results page follows which provides feedback on the outcomes of that round. The bid must be placed within $10$ seconds, otherwise the participant is signalled as inactive, and the bid defaults to zero. The results page communicates the outcome of the round in terms of whether priority is granted, the karma payment and redistribution, and the updated karma balance and game score. In addition, it gives feedback on the opposing bid (but not the opposing urgency and karma). Figures 45 in Appendix A show examples of these pages.
At the end of the game, participants are awarded a monetary payoff consisting of a fixed fee ${\phi_{\textup{fix}}}$ and a bonus fee ${\phi_{\textup{bon}}}$ that depends on the final score $s(T)$. Therefore, they are encouraged to be the winning bidder in as many rounds as possible, and especially in high urgency rounds. The bonus fee is determined according to the following rule:
  • If the participant is inactive for more than $6$ consecutive rounds, they are considered to not complete the experiment and are not awarded any payoff, i.e., ${\phi_{\textup{fix}}}={\phi_{\textup{bon}}}=0$;
  • Otherwise, the bonus fee is computed as an affine function of the score, given by
    ${\phi_{\textup{bon}}}$ $=\max\left\{({\phi_{\textup{targ}}}-{\phi_{\textup{rand}}})\>\frac{s(T)-{s_{\textup{rand}}}}{{s_{\textup{targ}}}-{s_{\textup{rand}}}}+{\phi_{\textup{rand}}},\>0\right\}.$
This rule linearly interpolates between two payment levels: ${\phi_{\textup{targ}}}$ is the payment associated with a target score ${s_{\textup{targ}}}$; and ${\phi_{\textup{rand}}}$ is the payment associated with the expected score for random bidding ${s_{\textup{rand}}}$. Hence, setting ${\phi_{\textup{rand}}}$ to a low value disincentivizes random play.
Participants were recruited on Amazon Mechanical Turk (MTurk). For each of the four treatment combinations, we ran five experiments with $N=20$ participants each, for a total of 100 participants per treatment and 400 participants overall. The fixed fee was ${\phi_{\textup{fix}}}=\$1.5$ and compensates for a maximum waiting time of 10 minutes to form an experiment group. The bonus fee ${\phi_{\textup{bon}}}$ was tuned based on the observed performance in technical pre-tests such that participants receive approximately ${\phi_{\textup{targ}}}=\$10$ on average, which compensates for a maximum experiment duration of 40 minutes.
Table 3 in Appendix A lists all parameter values used in the experiments.
3. Results
In order to present our findings, we must first introduce the welfare measures used to quantitatively assess our results.
3.1. Welfare Measures
Our central welfare measure is the efficiency gain which we define next. For a particular participant $i$, let $\left(u_{i}(t)\right)_{t\in\{1,\dots,T\}}$ be the vector of realized urgency in the experiment, $S_{i}=s_{i}(T)$ the total score at the end of the experiment, and $S^{\textup{rand}}_{i}=\frac{1}{2}\sum_{t=1}^{T}u_{i}(t)$ the ex-ante expected total score, given the urgency realization, under random allocation. Then the efficiency gain is defined as
(1) $E_{i}=\frac{S_{i}-S^{\textup{rand}}_{i}}{S^{\textup{rand}}_{i}},$
and gives the relative improvement of participant $i$ with respect to the ex-ante expected random allocation. On this basis, we will assess efficiency based on the median efficiency gain among participants, and fairness based on the distribution of efficiency gains.
3.2. High-level Synopsis of Results
Overall, we find that there are pronounced and statistically significant efficiency gains in all treatments, higher under the high stake process than under the low stake process, but not significantly different from one another in terms of overall efficiency gains depending on whether the bidding scheme is binary or full range. In all treatments, more than $90\%$ of the population is better off with karma than under random allocation. The most favorable combination is high stake with binary bidding, both in terms of median efficiency gains and in terms of the distribution of gains as most individuals are on the benefiting side.
Remark: With the best intentions, we had fully pre-registered design and analysis of our experiments. Some but not all of the analyses presented here were indeed pre-registered, and there are also further analyses from the pre-registration that are not presented in this paper. We stuck with the pre-registration as much as we could, but the unfortunate need to depart from the pre-registered plan of analysis is explained in Appendix B.
3.3. Efficiency Results
Efficiency went up in all treatments. Figure 2 shows the median efficiency gain in each of the four treatments (denoted ‘total’), cf. Table 1, as well the median efficiency gain in the bottom and top halves of the population in terms of individual outcomes. For the purpose of visualizing the data spread, estimates of the $95\%$ confidence intervals (CI) are also shown in the figure based on 1000 bootstraps. [^1] Note that data samples are dependent for participants in the same experiment, therefore the bootstrapped CIs should be seen as heuristic estimates and for visualization purposes. For the random allocations, the CI was estimated by running 1000 independent simulations.
[…]
Figure 2. Median efficiency gains for the four treatment combinations, cf. Table 1, with estimate of $95\%$ confidence interval. Also shown are the medians in the lower and upper halves of the population.
Table 2. Mann–Whitney–Wilcoxon (MWW) tests with test statistics $U$ (and associated $p$-values in brackets) of efficiency gains under karma versus random for all treatments. Karma treatments have $n_{\textup{karma}}=100$, and random is based on $n_{\textup{random}}=100\,000$ samples.
Richness of scheme
Binary Full range Combined
Urgency process
Low stake
6 187 810.06187810.06\,187\,810.06 187 810.0
($<0.001$)
6 591 766.56591766.56\,591\,766.56 591 766.5
($<0.001$)
25 557 373.025557373.025\,557\,373.025 557 373.0
($<0.001$)
High stake
6 640 154.56640154.56\,640\,154.56 640 154.5
($<0.001$)
6 242 150.06242150.06\,242\,150.06 242 150.0
($<0.001$)
25 765 437.025765437.025\,765\,437.025 765 437.0
($<0.001$)
Combined
25 701 703.025701703.025\,701\,703.025 701 703.0
($<0.001$)
25 566 897.025566897.025\,566\,897.025 566 897.0
($<0.001$)
102 525 023.5102525023.5102\,525\,023.5102 525 023.5
($<0.001$)
[…]
3.4. Fairness Results
Figure 2 shows that in addition to achieving overall efficiency gains, the karma allocation is more efficient than random allocation for both the lower/less fortunate and the higher/more fortunate halves of the population. To provide finer grained insight on the fairness of karma, Figure 3 further shows the mean efficiency gain per population decile for the four treatments. It is important to note that while the efficiency gain (1) is defined with respect to the ex-ante expected score under random allocation, not the whole population realizes this score ex-post. To account for the fact that there will be more or less fortunate individuals under random allocation, Figure 3 also shows the ex-post mean efficiency gains per population decile for random allocation, with the $95\%$ confidence interval estimated from 1000 independent simulations per treatment. The key feature to observe is that in all treatments, $90\%$ of the population achieve higher efficiency in the karma scheme than the random scheme. Only the lowest decile is worse-off in karma than random. This decile is dominated by non-adopters whose bids defaulted to zero due to inactivity (number of non-adopters: low stake-binary: 9/100; low stake-full range: 6/100; high stake-binary: 7/100; high stake-full range: 5/100). Non-adopters achieve particularly low scores as with a consistent bid of zero it is very unlikely to get granted priority.
Figure 3. Mean efficiency gain per decile for the four treatment combinations.
Moreover, it is evident that the treatment combination of high stake-binary is particularly favorable. Not only does it lead to the highest total median efficiency gain, cf. Figure 2, but it also achieves the largest gap to random allocation across deciles, cf. right panel of Figure 3. This suggests that all (active) subgroups of the population achieve strong benefits in this treatment.
Notice that the variability in the ex-post efficiency gains under random allocation, cf. Figure 3, can be improved by adopting simple turn-taking schemes, such as a fixed-value token scheme granting priority to participants holding the highest number of tokens (or equivalently, the lowest number of previous allocations). Such a scheme would achieve the same aggregate efficiency as random allocation, but ensure the ex-post efficiency gains are more equally distributed around zero in Figure 3. Therefore, in comparison to a simple turn-taking scheme, karma would also achieve pronounced efficiency gains overall, meanwhile the proportion of the population benefiting would drop to $70\%$ to $80\%$ depending on the treatment.
4. Discussion
In sum, in this paper we find that the aggregate efficiency gains of a karma scheme compared with random allocation are pronounced and statistically significant in all treatments. This constitutes the first set of behavioral evidence that a formal karma mechanism indeed can work to the benefit of the population. Moreover, in all treatments, (almost) all participants manage to benefit from the karma scheme in comparison to random allocation, with the exception of the lowest decile of non-adopters who did not actively participate in the bidding. Thus, our experiments provide the first set of formal evidence for the potential social benefits of using a karma scheme with human participants, as it improves efficiency to the benefit of almost everybody.
[…]
Another way to interpret the observed differences between online subjects and Nash or expert play may be related to impatience or perception of future rounds as less important. To investigate this, we compared the aggregate efficiency gains realized in our online experiments to those that would be achieved in theory in a short-sighted equilibrium, as predicted by (Elokda et al., 2023). It was found that the realized gains of online subjects correspond to if players discount future rounds by a factor of approximately 0.3–0.7, while those of experts correspond to a future discount factor of more than 0.98.
Another important consideration regarding implementation of karma with human participants is whether a simpler scheme (binary) or a richer scheme (full range) is favourable. That fact that we found no significant differences in terms of realized efficiency gains between the two provides preliminary evidence that the simpler binary scheme–at least for applications similar to the ones we studied–can be employed without loss in efficiency. Such applications feature, as we have investigated in our experiments, urgency processes with binary levels, for which a binary scheme is arguably also particularly natural. Theoretically, Nash equilibrium under binary bidding will lose in efficiency with more than two urgency levels, and it remains an open question to test what the trade-off is between behavioral simplicity of a binary (or otherwise limited) bidding scheme and the theoretical benefits that come with richer schemes. Some reasons to believe simplicity is beneficial are based on theories of decision fatigue (Baumeister, 2003; Pignatiello et al., 2020) and simplicity in mechanism design (Pycia and Troyan, 2023). A richer analysis of this complexity-performance trade-off provides fruitful avenues for future investigations, both in theory and in the behavioral lab.
Finally, to conclude, we would like to highlight that the most favorable treatment combination in our experiments was that of high stake urgency process under binary bidding scheme; both median efficiency gains and the distribution of gains were higher in that combination than in all others. Having rare but important urgency realizations made it particularly easy for our subjects to decide whether to bid or not, and there were no subtleties regarding how much to bid as a function of the history of play, etc. given the binary nature of the bidding scheme.
Appendix A Supplementary Implementation Details
Figure 4 shows examples of the decision page implemented in the online experiment. Figure 5 shows an example of the results page. Table 3 lists the notation and detailed parameter values used in the experiments.
(a) High stake, binary treatment.
(b) Low stake, full range treatment.
Figure 4. Decision page examples.
Figure 5. Results page example.
Table 3. List of notation and parameter values
Parameter Description Low stake High stake
$N$ Number of participants 20
$T$ Number of rounds 50
${k_{\textup{init}}}$ Initial karma 9
${k_{\textup{max}}}$ Max karma 18
${u_{\textup{l}}}$ Low urgency level 1
${u_{\textup{h}}}$ High urgency level 5 9
$\mathbb{P}({u_{\textup{h}}})$ High urgency probability 0.5 0.25
${s_{\textup{targ}}}$ Target score 90 101.25
${s_{\textup{rand}}}$ Random score 37.5
${\phi_{\textup{targ}}}$ Target bonus fee \$10
${\phi_{\textup{rand}}}$ Random bonus fee \$1
${\phi_{\textup{fix}}}$ Fixed fee \$1.5
${T_{\textup{test}}}$ Number of test rounds 5
${T_{\textup{dec}}}$ Decision inactivity timer 10s
${T_{\textup{inactive}}}$ Inactivity counter 6
Appendix B Remarks regarding departure from pre-registered analysis plan
Originally, we pre-registered the design of the experiment along with a complete plan for analysis at the Open Science Framework under [link withheld to preserve anonymity of submission]. The analysis presented in the paper above constitutes a deviation from this pre-registered analysis plan. Unfortunately, we had to re-design our analysis and statistical tests due to a technical bug in the logging of a participant’s bid who made a non-zero bid but was outbid (the ‘loser’ that round) by the other participant (the ‘winner’ that round), and therefore did not spend any karma. In scenarios like this, the winner and the karma accounting were done correctly, but a bug in the data-logging implementation resulted in a loss of data concerning the bid levels of the losers, thus making them indistinguishable also from players who bid zero. This loss of data logging made it impossible to perform the pre-registered analysis, and we opted for the alternative analysis presented above. Here, for completeness, we state the hypotheses that were originally pre-registered:
  1. (1)
    The karma allocation is more efficient than a random allocation.
  2. (2)
    The karma allocation is more fair than a random allocation.
  3. (3)
    The efficiency of the karma allocation is within 10% of the most efficient allocation.
  4. (4)
    The karma allocation is fairer than an efficiency-maximizing but history-unaware allocation.
  5. (5)
    There is a positive correlation between the participants’ bid and urgency.
  6. (6)
    A full bidding scheme is more efficient than a binary bidding scheme.
  7. (7)
    A binary bidding scheme is more fair than a full bidding scheme.
  8. (8)
    Participants with high urgency spread achieve higher rewards than those with low urgency spread.
As regards the efficiency-related hypotheses (1), (3), (6) and (8), the results of our MWW tests strongly support hypothesis (1), weakly support hypothesis (8), and reject hypotheses (3) and (6). Note, however, that our original plan was to measure realized efficiencies using ex-post means, which was found to be overly sensitive to stochastic effects. Therefore, the analysis presented in Section 3 is based on ex-ante expected medians instead. As regards the fairness-related hypotheses (2), (4) and (7), our fairness analysis supports hypotheses (2) and (7), in particular for the high stake treatments, but not hypothesis (4). These conclusions are drawn based on comparisons of the efficiency gains across deciles, and not as measured by the negated standard deviations of final scores, as we had originally planned, which turned out to be a measure that does not produce interpretable results. Unfortunately, due to the bug, hypothesis (5) could not be tested.
Appendix C Supplementary Results
Table 4 reports the MWW test results when comparing the efficiency gains of the first versus the second halves of the experiments, per treatment. With the exception of the high stake-binary treatment, no statistically significant differences are observed.
Table 4. Mann–Whitney–Wilcoxon (MWW) tests with test statistics $U$ (and associated $p$-values in brackets) of efficiency gains in the first versus second halves of the experiments for all treatments. All treatments have $n_{\textup{karma}}=100$ samples.
Richness of scheme
Binary Full range Combined
Urgency process
Low stake
5 005.05005.05\,005.05 005.0
($0.9912$)
5 460.55460.55\,460.55 460.5
($0.2610$)
20 878.520878.520\,878.520 878.5
($0.4476$)
High stake
4 200.54200.54\,200.54 200.5
($0.0509$)
4 949.54949.54\,949.54 949.5
($0.9028$)
18 357.518357.518\,357.518 357.5
($0.1555$)
Combined
18 238.518238.518\,238.518 238.5
($0.1277$)
20 684.020684.020\,684.020 684.0
($0.5544$)
77 729.077729.077\,729.077 729.0
($0.4872$)
Table 5 reports the MWW test results when comparing the efficiency gains of each treatment pair as well as between combined treatments. With the exception of low stake-binary versus high stake-binary (and consequently combined low stake versus high stake), no statistically significant differences are observed.
Table 5. Inter-treatment efficiency gain MWW tests reporting test statistics $U$ (with associated $p$-values in brackets). All treatments have $n=100$ samples.
Treatment comparisons
Low stake High stake Combined
Binary
Full range
Combined
Binary
Full range
Combined Binary Full range
Low stake
Binary
4 700.54700.54\,700.54 700.5
($0.4650$)
4 186.04186.04\,186.04 186.0
($0.0468$)
4 522.04522.04\,522.04 522.0
($0.2433$)
Full range
5 299.55299.55\,299.55 299.5
($0.4650$)
4 412.04412.04\,412.04 412.0
($0.1511$)
4 744.04744.04\,744.04 744.0
($0.5324$)
Combined
17 864.017864.017\,864.017 864.0
($0.0647$)
High stake
Binary
5 814.05814.05\,814.05 814.0
($0.0468$)
5 588.05588.05\,588.05 588.0
($0.1511$)
5 288.05288.05\,288.05 288.0
($0.4824$)
Full range
5 478.05478.05\,478.05 478.0
($0.2433$)
5 256.05256.05\,256.05 256.0
($0.5324$)
4 712.04712.04\,712.04 712.0
($0.4824$)
Combined
17 864.017864.017\,864.017 864.0
($0.0647$)
Combined
Binary
20 098.520098.520\,098.520 098.5
($0.9324$)
Full range
20 098.520098.520\,098.520 098.5
($0.9324$)
Figure 6 compares the median efficiency gains realized in our online MTurk experiments with a) the gains that would have been achieved if the participants followed the Nash equilibrium (as determined by running 1000 simulations), and b) for the low stake-full range treatment, the gains achieved by ‘expert’ subjects, those are, students of an applied game theory class [^4] The median efficiency gain of Expert is higher than Nash due to a large number of inactive participants (9/28). Inactive participants were added due to a technical issue in forming an experimental group that required setting the size of the group greater than the number of students.. This suggests that stronger benefits can be achieved with more training and commitment.
Figure 6. Comparison of the median efficiency gains in the online MTurk experiments versus simulations of the Nash equilibrium. For the low stake-binary treatment, the median efficiency gain from an ‘expert’ game theory class experiment is also shown.