### v2

# Dynamic Recourse Allocation with Karma: An Experimental Study

Abstract
A system of non-tradable credits that flow between individuals like karma, hence proposed under that name, is a mechanism for repeated resource allocation that comes with attractive efficiency and fairness properties, in theory. In this study, we test karma in an online experiment in which human subjects repeatedly compete for a resource with time-varying and stochastic individual preferences or urgency to acquire the resource. We confirm that karma has significant and sustained welfare benefits even in a population with no prior training. We identify mechanism usage in contexts with sporadic high urgency, more so than with frequent moderate urgency, and implemented as a simple (binary) karma bidding scheme as particularly effective for welfare improvements: relatively larger aggregate efficiency gains are realized that are (almost) Pareto superior. These findings provide guidance for further testing and for future implementation plans of such mechanisms in the real world.
keywords:
Behavioral economics , Repeated allocation , Karma economy , Artificial currency
journal: Games and Economic Behavior
[…]
[ifa] organization=Automatic Contol Lab, ETH Zürich,addressline=Physikstrasse 3, city=Zürich, postcode=8092, country=Switzerland
[…]
[uzh] organization=Department of Sociology, University of Zürich,addressline=Andreastrasse 15, city=Zürich, postcode=8050, country=Switzerland
1 Introduction
Figure 1: Dynamic resource allocation with karma involving three cars and a repeated sequence of three encounters. Start with the encounter in the top center. The high-urgency lila car has a current karma account of 9 and bids 6, thus outbidding the low-urgency blue car whose karma account is also 9 but bids 2. The bids of the lila car (and, in general, of other cars winning in parallel encounters) are paid and redistributed. As a result, the blue car’s karma account goes up by 3 to 12. Let’s move along clockwise. Blue now happens to have high urgency and bids 6, thus outbidding and getting priority over orange whose karma account is 9 and bids 3. Now the orange karma goes up by the redistribution share of 3 to 12. Let’s move along clockwise. Orange now has high urgency and bids 6, thus outbidding low-urgency lila who has 6 karma left and bids 1. Thus the circle closes, et cetera, et cetera.
Efficiency and fairness in determining who gets what and when are the two major objectives in resource allocation situations under scarcity, and many interesting mechanism and market design solutions have been proposed (Roth, 2015). An important class of allocation problems is when goods are repeatedly and indefinitely allocated amongst a fixed population: for example, farmers require daily access to shared groundwater resources, commuters require regular access to roads, students require frequent access to scarce computing clusters, food banks require daily access to food donations, etc. In situations like these, it is often the case that one person relative to another gets higher utility from the good in one period but lower utility in another, which we shall refer to as time-varying levels of urgency. A recent innovation to address exactly those kinds of allocation problems goes under the name karma (Vishnumurthy et al., 2003; Elokda et al., 2023; Vuppalapati et al., 2023). The karma mechanism mirrors what Western popular culture associates with the notions of karma and samsara stemming from Indian religions according to which one’s deeds in the present (karma) affect the quality of one’s future life (phala) and thereafter (samsara) (Reichenbach, 1990; Kyabgon, 2015). The mechanism proposed in that literature is implemented via individual accounts of non-tradable credits called karma. Individuals may bid some amount from their current account at every instance of resource allocation. In the baseline implementation of the mechanism, the highest bidders win and obtain priority for the resource, and must pay their bids, which are then redistributed in the population [^1] This corresponds to the first-price auction implementation of the mechanism. Other auctions, like second-price, have also been discussed.. Figure 1 illustrates how the karma mechanism works out to the benefit of everyone at hand of an example involving three infinitely re-occurring road intersection encounters.
The karma mechanism is simple and appealing, but not yet frequently used in practice. An exception is the ‘choice system’ proposed by Prendergast (2022), which is adopted by U.S. food banks, where karma tokens are called ‘shares.’ Food banks under the choice system bid shares to obtain priority for available food donations, pay their bids upon winning based on a first-price auction mechanism, and the total payment of shares is redistributed at the end of each day [^2] Indeed, Prendergast (2022) discusses alternatives to first-price auctions and their potential benefits, but mentions that the real-world implementation partners had a strong preference for first-price owed to its transparency and simplicity.. This real-world example is emblematic of the kind of resource allocations for which karma is suited: the use of monetary transfers is deemed inappropriate or highly undesirable; the allocations are repeated frequently (i.e., daily); and, importantly, there is no finite horizon in sight for when the allocations will cease to exist. Indeed, the design choice of redistributing the paid shares is critical to address the infinite repetition of the allocation: by forming a ‘closed economy’ in which total shares are preserved over time, a stationary regime can be reached and henceforth repeated indefinitely.
The idea of sacrificing resource consumption today in favor of future consumption in periods of higher urgency is intuitive, and it is not surprising that this same idea underpins several related lines of works, including linking decisions across periods (Jackson and Sonnenschein, 2007; Hortala-Vallve, 2010; Escobar and Toikka, 2013), trading favors (Möbius, 2001; Olszewski and Safronov, 2018b, a; Leo, 2017), token or artificial currency-based mechanisms (Johnson et al., 2014; Gorokh et al., 2021a, b; Banerjee et al., 2023), and trading votes (Casella and Macé, 2021; Casella, 2005; Casella et al., 2006; Hortala-Vallve and Llorente-Saguer, 2010; Hortala-Vallve, 2012; Casella and Palfrey, 2019, 2021; Macías, 2024). With respect to these works, which are discussed in Section 2, karma and the aforementioned choice system have the distinguishing feature of forming closed economies that are particularly suited for infinitely repeated resource allocations.
The choice system, which has indeed resulted in significant aggregate gains in terms of food bank participation and fluidity of donations (Prendergast, 2022), provides empirical evidence that karma mechanisms can be successful in practice. On the theory side, recent works have studied game-theoretical models of karma mechanisms in simplified, single repeated resource allocation settings (Elokda et al., 2023, 2024). These works have been motivated by the use of karma as a public policy instrument that is an equitable alternative to classical (monetary) congestion pricing policies, e.g., for allocating priority roads and other public infrastructures. However, in order to realize the potential benefits of karma-based policies, and employ these policies effectively in human populations, systematic behavioral evidence is needed. The present paper contributes to this emergent strand of literature a first controlled experimental test of karma. We conduct an experiment on karma with different bidding schemes and different urgency processes in order to better understand the behavioral uptake of karma by human actors and its efficiency and distributional consequences.
The theoretical predictions of Elokda et al. (2023, 2024) are based on the solution concept of the Stationary Nash Equilibrium (SNE): a compact, time-invariant predictor of optimal rational behavior, which is guaranteed to exist due to the preservation of karma. Elokda et al. (2023) develop computation tools for the SNE and show that farsighted SNE are almost fully efficient with respect to the private urgency of the users. These predictions are based on idealized assumptions including rationality, far-sightedness, infinite population, perfect adoption, and that the stationary conditions for which the SNE is optimal are reached. In this paper, we perform a behavioral investigation with real humans who do not necessarily follow such idealized assumptions. We want to know whether humans find it natural to adopt the mechanism, whether inexperienced players are able to realize efficiency gains, and whether stationarity is attained in practice. In our analysis we focus on the overall efficiency effects of karma as well as on distributional consequences and associated fairness properties of the mechanism. In our treatments, we vary the nature of the dynamic urgency distributions (more frequent and less intense versus less frequent and more intense) and the auction process of the karma mechanism (binary bidding versus full bidding). We use random allocation as a benchmark to compare our findings to. This domain-independent benchmark is representative of common schemes that are unaware of the individual private urgency, including fixed turn-taking schemes and schemes using fixed-value tokens.
The main insights of our experiments summarize as follows:
  • 1.
    Almost all participants benefit in the karma allocation compared with random allocation: this is the case for 90% of the participants, and the remaining 10% are mostly non-adopters who do not participate in karma bidding actively themselves. If we look at active participants only, karma led to an almost Pareto improvement.
  • 2.
    The Pareto improvement occurs despite the relatively low level of training and commitment by the online participants.
  • 3.
    The realized benefits, while greater than in random allocation, fall short of theoretical Nash predictions. Analysis of the bidding behaviors reveals that the main deviation to Nash behavior takes the form of irrational over-bidding in low urgency rounds.
  • 4.
    Benefits are particularly pronounced in situations when preference intensities are dynamically more intense and less frequent, and the bidding scheme is designed to be minimal (i.e., binary).
  • 5.
    The realized variations in the karma and bid distributions over time are small and comparable in magnitude to the variations that would be attained if all participants followed stationary Nash behavior. This suggests that an approximately stationary regime is reached despite the presence of noisy bidding behaviors. Moreover, the variations are smaller under binary bidding than full bidding, suggesting that the karma auctions are particularly predictable under binary bidding.
These findings provide a first benchmark that karma may be used beneficially and robustly in human interactions. Our study also points in several directions for further theoretical and experimental investigation.
2 Related Mechanisms
In this section, we highlight how karma mechanisms differ from several previously proposed mechanisms that share the same intuitive idea of trading off between present or future access to resources.
Linking decisions
Mechanisms based on linking decisions (Jackson and Sonnenschein, 2007; Hortala-Vallve, 2010; Escobar and Toikka, 2013) rely on correlating each individual’s reports over time with publicly known distributions of the private urgency, and punishing those individuals that deviate. This requires to keep track of individual identities and histories, and we view karma and other tokens as memory-efficient instruments to link decisions that are capable of scaling in large populations.
Trading favors
Mechanisms based on trading favors (Möbius, 2001; Olszewski and Safronov, 2018b) rely on simple book-keeping of favors owed, but these classical mechanisms are tailored to truthful reporting of one’s availability to grant favors with no regard to time-varying private urgency. Leo (2017) addresses time-varying urgency specifically for two individuals taking turns to perform chores, while Olszewski and Safronov (2018a) addresses time-varying urgency in more general settings using probabilistic exchange of karma-like tokens called ‘chips’. However, Olszewski and Safronov (2018a)’s mechanism depends on the individual preference distributions in a complex manner and thus does not scale naturally.
Tokens and artificial currency
Karma-like instruments have been previously referred to as vouchers, tokens, scrips, or artificial currency. We distinguish between token-based mechanisms in which the value of the resource is fixed in tokens (Johnson et al., 2014) (typically one resource unit is worth one token); and artificial currency-based mechanisms in which, like karma, the value of the resource is determined in an auction-like mechanism (Gorokh et al., 2021a, b; Banerjee et al., 2023). Token-based mechanisms are not well suited to elicit time-varying private urgency; whereas most previously proposed artificial currency-based mechanisms are tailored to finite resource repetitions: individuals are issued an initial budget of currency to spend over the finite horizon (with no redistribution or other forms of currency exchange). The non-preservation of total system currency makes these mechanisms less well-suited to infinite resource repetitions: they would require a periodic central endowment of currency (e.g., every month or year), do not forgive mistakes leading to early depletion of currency, and, importantly, lead to non-stationary settings in which optimal strategies depend explicitly on the time left in the horizon.
Trading votes
Trading votes across issues or proposals is an intuitively appealing and practically prevalent practice, yet it remains unclear to what extent vote trading improves welfare and how to design vote trading mechanisms optimally, as pointed out in Casella and Macé (2021)’s recent review. Casella and Macé (2021) distinguish between two types of vote trading: those in which votes are traded with other voters (Casella and Palfrey, 2019, 2021); and those in which votes are traded individually with one’s future self (referred to as storable (Casella, 2005) or qualitative votes (Hortala-Vallve, 2012)). The latter type of storable votes, which yields particularly favorable efficiency gains in comparison to the other types of vote trading (Casella and Macé, 2021), is closely related to the aforementioned class of (finite-horizon) artificial currency mechanisms: voters are issued an initial budget of votes to cast in a (small) finite number of issues (Casella, 2005; Casella et al., 2006; Hortala-Vallve, 2012). One recently proposed mechanism by Macías (2024) resembles karma more closely: in this mechanism, votes are “paid” by the majority voters and subsequently redistributed, however, Macías (2024) studies a two voter only model. Our study thus complements the literature on storable votes, as we are motivated by resource allocations that are repeated more frequently and typically involve more players than in voting. In contrast to previous experimental studies on vote trading (Casella et al., 2006; Hortala-Vallve and Llorente-Saguer, 2010; Casella and Palfrey, 2019), our experiments involve significantly more rounds and larger groups.
3 Experimental Methods
We conducted a balanced two-by-two factorial experiment with 400 subjects in total. Treatments varied in the dynamic urgency process of the participants and the richness of the karma scheme. For urgency, we distinguished between a low stake process where participants have frequent events with moderate urgency, and a high stake process where participants have rare events with high urgency. For richness, we tested a binary scheme where participants can choose from only two bid levels that depend on their karma, and a full range scheme where participants have full choice over the bid up to their karma.
3.1 The Game
The game we study is one that proceeds with $N$ participants over $T$-many rounds. All participants receive an initial endowment of karma $k(0)=k(1)={k_{\textup{init}}}\in\mathbb{N}$ and an initial game score $s(0)=0$. At each round $t\in\{1,\dots,T\}$, participants are randomly matched in pairs to compete over a shared resource. Each participant is given an urgency value $u(t)\in\{{u_{\textup{l}}},\>{u_{\textup{h}}}\}$ that is drawn randomly and independently from a process $\mathbb{P}(u)$ that is identical for all players, and must place a bid $b(t)\leq k(t)$. In each pairwise matching, the higher bidder gets allocated the resource: the score $s(t)=s(t-1)+u(t)$ increases by the urgency and the bid is collected as payment. The lower bidder does not get allocated the resource: the score $s(t)=s(t-1)$ does not increase and no payment is collected. Ties are settled randomly. To keep the total amount of karma constant, at the end of each round, the total collected payment ${p_{\textup{tot}}}$ is uniformly redistributed to all participants in an integer preserving manner. In case ${p_{\textup{tot}}}$ is not divisible by $N$, $\left\lfloor\frac{{p_{\textup{tot}}}}{N}\right\rfloor$ is redistributed to everyone, and a random subset of ${p_{\textup{tot}}}-N\left\lfloor\frac{{p_{\textup{tot}}}}{N}\right\rfloor$ participants receive one additional karma. Moreover, each participant is allowed a maximum karma level ${k_{\textup{max}}}$ that the redistribution respects. Participants with ${k_{\textup{max}}}$ do not receive additional redistribution which instead gets issued uniformly to the others. After the redistribution the game proceeds to the next round $t+1$.
Table 3 in A summarizes the notation introduced above.
3.2 Treatments
We follow a two-by-two factorial treatment design, where we vary the dynamic urgency process of the participants and the richness of the karma scheme. Table 1 summarizes the treatment configurations schematically. For each of the four resulting treatments we run 5 independent game experiments with 20 participants per experimental game resulting in 100 subjects per treatment and 400 subjects in total.
Table 1: 2x2 experimental design.
Richness of scheme
Binary Full range
Urgency process
Low stake [Uncaptioned image]    [Uncaptioned image] [Uncaptioned image]    [Uncaptioned image]
High stake [Uncaptioned image]    [Uncaptioned image] [Uncaptioned image]    [Uncaptioned image]
3.2.1 Urgency Process
The urgency processes in both treatment variations of low stake and high stake have the same low urgency ${u_{\textup{l}}}=1$, but differ in the magnitude and frequency of the high urgency (low stake: ${u_{\textup{h}}}=5$, ${\mathbb{P}}({u_{\textup{h}}})=0.5$; high stake: ${u_{\textup{h}}}=9$, ${\mathbb{P}}({u_{\textup{h}}})=0.25$). The motivation for these treatment variations is to investigate behavioral effects under urgency processes of different dynamic nature. Both processes have the same urgency on average $\mathbb{E}(u)=3$, and therefore the same expected scores under random allocation. Low stake represents the case where the high urgency event is frequent but moderate, while in high stake the high urgency event is rare but severe. Notice that due to the different dynamic nature of the processes, there is a greater potential to benefit over random allocation in the high stake process than in the low stake process (cf. Nash efficiency gains in Figure 2).
3.2.2 Richness of Karma Scheme
In the treatment variation of binary, participants can only choose between two bid levels: 00 or $\left\lfloor\frac{k}{2}\right\rfloor$ (i.e., half their karma, rounded down to the nearest integer). In the treatment variation of full range, participants can choose any integer bid up to their karma. The motivation of these treatment variations is to investigate the behavioral effects of a reduced action space, i.e., binary, where subjects either bid or not. The particular design choice to restrict bids to 00 or $\left\lfloor\frac{k}{2}\right\rfloor$ in the binary scheme was guided by theoretical Nash predictions: using tools from Elokda et al. (2023), these binary bid levels were predicted to achieve almost the same efficiency at the Nash equilibrium as in full range bidding, for the urgency processes considered (cf. Nash efficiency gains in Figure 2). Therefore, the simpler binary scheme does not trade-off performance, in theory, and any differences observed in the outcomes of the two bidding schemes will be due to behavioral effects.
3.3 Experimental Implementation
The game is implemented as a real-time online experiment using oTree (Chen et al., 2016). Participants are given the opportunity to familiarize with the game over $5$ rounds that do no contribute to the game score, after which the main game proceeds with $T=50$ rounds. In each round, participants get drawn a random urgency. On the decision page, they must place a bid, and then a results page follows which provides feedback on the outcomes of that round. The bid must be placed within $10$ seconds, otherwise the participant is signalled as inactive, and the bid defaults to zero. The results page communicates the outcome of the round in terms of whether priority is granted, the karma payment and redistribution, and the updated karma balance and game score. In addition, it gives feedback on the opposing bid (but not the opposing urgency and karma). Figures 78 in A show examples of these pages.
At the end of the game, participants are awarded a monetary payoff consisting of a fixed fee ${\phi_{\textup{fix}}}$ and a bonus fee ${\phi_{\textup{bon}}}$ that depends on the final score $s(T)$. Therefore, they are encouraged to be the winning bidder in as many rounds as possible, and especially in high urgency rounds. The bonus fee is determined according to the following rule:
  • 1.
    If the participant is inactive for more than $6$ consecutive rounds, they are considered to not complete the experiment and are not awarded any payoff, i.e., ${\phi_{\textup{fix}}}={\phi_{\textup{bon}}}=0$;
  • 2.
    Otherwise, the bonus fee is computed as an affine function of the score, given by
    ${\phi_{\textup{bon}}}$ $=\max\left\{({\phi_{\textup{targ}}}-{\phi_{\textup{rand}}})\>\frac{s(T)-{s_{\textup{rand}}}}{{s_{\textup{targ}}}-{s_{\textup{rand}}}}+{\phi_{\textup{rand}}},\>0\right\}.$
This rule linearly interpolates between two payment levels: ${\phi_{\textup{targ}}}$ is the payment associated with a target score ${s_{\textup{targ}}}$; and ${\phi_{\textup{rand}}}$ is the payment associated with the expected score for random bidding ${s_{\textup{rand}}}$. Hence, setting ${\phi_{\textup{rand}}}$ to a low value disincentivizes random play.
Participants were recruited on Amazon Mechanical Turk (MTurk). For each of the four treatment combinations, we ran five experiments with $N=20$ participants each, for a total of 100 participants per treatment and 400 participants overall. The fixed fee was ${\phi_{\textup{fix}}}=\$1.5$ and compensates for a maximum waiting time of 10 minutes to form an experiment group. The bonus fee ${\phi_{\textup{bon}}}$ was tuned based on the observed performance in technical pre-tests such that participants receive approximately ${\phi_{\textup{targ}}}=\$10$ on average, which compensates for a maximum experiment duration of 40 minutes.
Table 3 in A lists all parameter values used in the experiments.
4 Results
In order to present our findings, we must first introduce the welfare measures used to quantitatively assess our results.
4.1 Welfare Measures
Our central welfare measure is the efficiency gain which we define next. For a particular participant $i$, let $\left(u_{i}(t)\right)_{t\in\{1,\dots,T\}}$ be the vector of realized urgency in the experiment, $S_{i}=s_{i}(T)$ the total score at the end of the experiment, and $S^{\textup{rand}}_{i}=\frac{1}{2}\sum_{t=1}^{T}u_{i}(t)$ the expected total score, given the urgency realization, under random allocation. Then the efficiency gain of participant $i$ is defined as
$E_{i}=\frac{S_{i}-S^{\textup{rand}}_{i}}{S^{\textup{rand}}_{i}}.$ (1)
This definition expresses the relative improvement with respect to the expected random score given the urgency realization, in order to control for randomness in the urgency realization. On this basis, we will assess overall efficiency based on the median efficiency gain among participants, and fairness based on the distribution of efficiency gains.
[…]
4.2 High-level Synopsis of Results
Overall, we find that there are pronounced and statistically significant efficiency gains in all treatments, higher under the high stake process than under the low stake process, but not significantly different from one another in terms of overall efficiency gains depending on whether the bidding scheme is binary or full range. In all treatments, more than $90\%$ of the population is better off with karma than under random allocation. The most favorable combination is high stake with binary bidding, both in terms of median efficiency gains and in terms of the distribution of gains as most individuals achieve pronounced benefits.
Remark: With the best intentions, we had fully pre-registered design and analysis of our experiments. Some but not all of the analyses presented here were indeed pre-registered, and there are also further analyses from the pre-registration that are not presented in this paper. We stuck with the pre-registration as much as we could, but the unfortunate need to depart from the pre-registered plan of analysis is explained in B.
4.3 Efficiency Results
Efficiency went up in all treatments. Figure 2 shows the median efficiency gain in each of the four treatments (denoted “Total”), cf. Table 1, as well the median efficiency gain in the bottom and top halves of the population in terms of individual outcomes, contrasted to the median gains that would be realized if all participants followed theoretically optimal Nash behavior. For the purpose of visualizing the data spread, estimates of the $95\%$ confidence intervals (CI) are also shown in the figure based on 1000 bootstraps [^3] Note that data samples are dependent for participants in the same experiment, therefore the bootstrapped CIs should be seen as heuristic estimates for visualization purposes. Similar estimates were attained for larger numbers of bootstraps.. For the random and Nash allocations, the CI was estimated by running 1000 independent simulations.
The findings of Figure 2 summarize as follows. Although the experimental efficiency gains fall short of the theoretical Nash gains, there are nonetheless positive total efficiency gains in all treatments (median ranging from $7.38\%$ to $15.34\%$). The efficiency gains are especially pronounced in the upper half of the population (median ranging from $22.51\%$ to $29.85\%$); whereas the lower half of the population does not perform much worse than ex-ante expected under random allocation (median ranging from $-6.5\%$ to $-0.65\%$), and in fact performs better than the lower half of the population under random allocation ex-post. This suggests that overall, karma provides an opportunity for participants to achieve pronounced benefits, without strongly harming those that are less strategic. Moreover, the efficiency gains are statistically significant in all treatments, as elaborated in Table 2 reporting MWW test statistics $U$ (with associated $p$-values).
Figure 2: Median efficiency gains for the four treatment combinations, cf. Table 1, with estimate of $95\%$ confidence interval. Also shown are the medians in the lower and upper halves of the population.
Table 2: Mann–Whitney–Wilcoxon (MWW) tests with test statistics $U$ (and associated $p$-values in brackets) of efficiency gains under karma versus random for all treatments. Karma treatments have $n_{\textup{karma}}=100$, and random is based on $n_{\textup{random}}=100\,000$ samples.
Richness of scheme
Binary Full range Combined
Urgency process
Low stake
6 187 810.06187810.06\,187\,810.06 187 810.0
($<0.001$)
6 591 766.56591766.56\,591\,766.56 591 766.5
($<0.001$)
25 557 373.025557373.025\,557\,373.025 557 373.0
($<0.001$)
High stake
6 640 154.56640154.56\,640\,154.56 640 154.5
($<0.001$)
6 242 150.06242150.06\,242\,150.06 242 150.0
($<0.001$)
25 765 437.025765437.025\,765\,437.025 765 437.0
($<0.001$)
Combined
25 701 703.025701703.025\,701\,703.025 701 703.0
($<0.001$)
25 566 897.025566897.025\,566\,897.025 566 897.0
($<0.001$)
102 525 023.5102525023.5102\,525\,023.5102 525 023.5
($<0.001$)
[…]
4.4 Fairness Results
Figure 2 shows that in addition to achieving overall efficiency gains, the karma allocation is more efficient than random allocation for both the lower/less fortunate and the upper/more fortunate halves of the population. To provide finer grained insight on the fairness of karma, Figure 3 further shows the mean efficiency gain per population decile for the four treatments. It is important to note that while the efficiency gain (1) is defined with respect to the ex-ante expected score under random allocation given the urgency realization, not the whole population realizes this score ex-post. To control for the fact that there will be more or less fortunate individuals due to radnomness under random allocation, Figure 3 also shows the ex-post mean efficiency gains per population decile for random allocation, with the $95\%$ confidence interval estimated from 1000 independent simulations per treatment. The key feature to observe is that in all treatments, $90\%$ of the population achieve higher efficiency in the karma scheme than the random scheme. Only the lowest decile is worse-off in karma than random. This decile is dominated by non-adopters whose bids defaulted to zero due to inactivity (number of non-adopters: low stake-binary: 9/100; low stake-full range: 6/100; high stake-binary: 7/100; high stake-full range: 5/100). Non-adopters achieve particularly low scores as with a consistent bid of zero it is very unlikely to get granted priority.
Figure 3: Mean efficiency gain per decile for the four treatment combinations.
Moreover, it is evident that the treatment combination of high stake-binary is particularly favorable. Not only does it lead to the highest total median efficiency gain, cf. Figure 2, but it also achieves the largest gap to random allocation across deciles, cf. right panel of Figure 3. This suggests that all (active) subgroups of the population achieve pronounced benefits in this treatment.
Notice that the variability in the ex-post efficiency gains under random allocation, cf. Figure 3, can be improved by adopting simple turn-taking schemes, such as a fixed-value token scheme granting priority to participants holding the highest number of tokens (or equivalently, the lowest number of previous allocations). Such a scheme would achieve the same aggregate efficiency as random allocation, but ensure the ex-post efficiency gains are more equally distributed around zero in Figure 3. Therefore, in comparison to a simple turn-taking scheme, karma would also achieve pronounced efficiency gains overall, meanwhile the proportion of the population benefiting would drop to $70\%$ to $80\%$ depending on the treatment.
4.5 Analysis of Bidding Behaviors
[…]
(a) Low stake, Binary.
(b) Low stake, Full range.
(c) High stake, Binary.
(d) High stake, Full range.
Figure 4: Median bid per urgency and karma for the four treatment combinations.
The main finding of Figure 4 is that overall, there was a tendency to over-bid in the low urgency state [^4] Notice that due to the aforementioned bug which excluded some of the losing bids, the medians in Figure 4 are potentially biased towards higher winning bids. Nonetheless, there are sufficiently many samples to support the observed tendency to over-bid in low urgency.. The bidding behaviors are not well-explained by a single value of future discounting: the median bids are well-fitted to far-sighted behavior (discount factors 0.80.80.80.80.980.980.980.98) in high urgency, but to short-sighted behavior (discount factor $0.6$) in low urgency. In contrast, the median bids of top-performing participants are well-fitted to far-sighted behavior in both urgency levels, which explains the high efficiency gains achieved by these participants.
To provide further insight on how over/under-bidding during low/high urgency affected performance, Figure 5 shows a scatter of the mean signed difference to Nash-$0.98$ bids per participant and urgency level, versus the attained efficiency gains, for the four treatments. The data points associated to the top-performing decile are highlighted in orange. As expected, under-bidding in high urgency negatively affects performance, cf. third quadrant of the high urgency panels, which includes the majority of participants achieving negative efficiency gains. On the other hand, the effect of over-bidding in either of the urgency levels is inconclusive. Consistently in all treatments, the single-top performing participant bid close to Nash [^5] The mean absolute difference to Nash-$0.98$ bids correlates well with the efficiency gain, as expected, and particularly so in the full-range treatments (Spearman correlations: low stake, binary: $-0.345$, $p=0.0004$; low stake, full-range: $-0.664$, $p=0.0000$; high stake, binary: $-0.119$, $p=0.236$; high stake, full-range: $-0.674$, $p=0.0000$). However, several top performing participants also tended to over-bid in low urgency, and a few of these participants managed to over-bid in both low and high urgency, cf. Figure 5(d).
(a) Low stake, Binary.
(b) Low stake, Full range.
(c) High stake, Binary.
(d) High stake, Full range.
Figure 5: Efficiency gain vs. mean difference to Nash for the four treatment combinations.
4.6 Analysis of stationarity
An important feature of the karma mechanism is that it forms a closed economy in which karma is preserved over time. In theory, this feature enables reaching a stationary regime, in which a predictable, time-invariant optimal behavior can be repeated indefinitely. To investigate whether stationarity is achieved in practice, each sub-figure of Figure 6 shows the variations in the distribution of karma (left panel) and bids (right panel) over the course of the main experimental rounds, for each of the four treatments. The variation in the distributions is measured by the Wasserstein-1 distance, also known as the Earth mover’s distance, between the realized karma/bid distributions of successive rounds. The choice of the Wasserstein-1 distance is interpretable: for example, a distance of one corresponds to shifting the distribution by one karma unit on average. For each treatment, the realized distribution variations are plotted for each of the five experimental groups (labelled “MTurk-G1–G5”), and contrasted to the variations that would be attained if all participants followed stationary Nash behavior (labelled “Nash”; the shaded area coincides with the 95% confidence interval estimated from 1000 independent simulations per group).
(a) Low stake, Binary.
(b) Low stake, Full range.
(c) High stake, Binary.
(d) High stake, Full range.
Figure 6: Variation in the karma and bid distributions over time for the four treatment combinations.
The main finding of Figure 6 is that in all treatment combinations, after a short initial transient of less than five rounds, the realized variations in the karma and bid distributions are close in magnitude to those attained under stationary Nash play, as visualized by the shaded Nash area subsuming most of the experimental data. Two exceptions are the bid distributions in the full range treatments, cf. right panels of Figures 6(b), 6(d), in which the experimental variations lie at a notably higher magnitude than the stationary Nash variations. Nonetheless, these variations are still relatively small and oscillate around a Wasserstein-1 distance of one. Overall, this finding suggests that despite the presence of noisy bids, the karma auctions quickly become predictable over time, and these auctions are even more predictable in the binary treatments.
5 Discussion
In sum, in this paper we find that the aggregate efficiency gains of a karma scheme compared with random allocation are pronounced and statistically significant in all treatments. This constitutes the first set of behavioral evidence that a formal karma mechanism indeed can work to the benefit of the population. Moreover, in all treatments, (almost) all participants manage to benefit from the karma scheme in comparison to random allocation, with the exception of the lowest decile of non-adopters who did not actively participate in the bidding. Thus, our experiments provide the first set of formal evidence for the potential social benefits of using a karma scheme with human participants, as it improves efficiency to the benefit of almost everybody.
It is noteworthy that our experimental subjects were recruited from a population of totally untrained and inexperienced users from an online convenience sample (on MTurk). Their behavior, even though more efficient than random, however, is not as efficient as is theoretically feasible under Nash equilibrium play. Analysis of the experimental bidding behaviors reveals a consistent tendency to over-bid in low urgency rounds, however, this observation does not explain the gap to Nash efficiency entirely, which is likely attributed to irrational, noisy behavior. A natural follow-up question is, therefore, whether karma could be capable of achieving higher efficiency gains than realized in our online experiments if the human population consisted of participants that were better trained. As a first step in this direction, we conducted an auxiliary experiment under the low stake-full range treatment with a group of ‘expert’ subjects, who were graduate students in an applied game theory class. It turned out that, indeed, the achieved efficiency gains of these experts are close to Nash levels (MTurk: median $12.86\%$, $n=100$; experts: median $36.65\%$, $n=28$, MWW test $U=667.0$, $p=0.0000$). Therefore, we may interpret the efficiency gains that were realized in our online experiment as behaviorally robust lower bounds on the performance of the karma scheme given the relatively low training and commitment of the subject pool considered.
Another important consideration regarding implementation of karma with human participants is whether a simpler scheme (binary) or a richer scheme (full range) is favourable. That fact that binary led to less variations in the distribution of bids and thus more predictable auction outcomes, meanwhile we found no significant differences in terms of realized efficiency gains compared to full range, provides preliminary evidence that the simpler binary scheme is advantageous–at least for applications similar to the ones we studied. Such applications feature, as we have investigated in our experiments, urgency processes with binary levels, for which a binary scheme is arguably also particularly natural. Theoretically, Nash equilibrium under binary bidding will lose in efficiency with more than two urgency levels, and it remains an open question to test what the trade-off is between behavioral simplicity of a binary (or otherwise limited) bidding scheme and the theoretical benefits that come with richer schemes. Some additional reasons to believe simplicity is beneficial are based on theories of decision fatigue (Baumeister, 2003; Pignatiello et al., 2020) and simplicity in mechanism design (Pycia and Troyan, 2023). A richer analysis of this complexity-performance trade-off provides fruitful avenues for future investigations, both in theory and in the behavioral lab.
Finally, to conclude, we would like to highlight that the most favorable treatment combination in our experiments was that of high stake urgency process under binary bidding scheme; both median efficiency gains and the distribution of gains were higher in that combination than in all others. Having rare but important urgency realizations made it particularly easy for our subjects to decide whether to bid or not, and there were no subtleties regarding how much to bid as a function of the history of play, etc. given the binary nature of the bidding scheme.
Appendix A Supplementary Implementation Details
Figure 7 shows examples of the decision page implemented in the online experiment. Figure 8 shows an example of the results page. Table 3 lists the notation and detailed parameter values used in the experiments.
(a) High stake, binary treatment.
(b) Low stake, full range treatment.
Figure 7: Decision page examples.
Figure 8: Results page example.
Table 3: List of notation and parameter values
Parameter Description   Low stake   High stake
$N$ Number of participants 20
$T$ Number of rounds 50
${k_{\textup{init}}}$ Initial karma 9
${k_{\textup{max}}}$ Max karma 18
${u_{\textup{l}}}$ Low urgency level 1
${u_{\textup{h}}}$ High urgency level 5 9
$\mathbb{P}({u_{\textup{h}}})$ High urgency probability 0.5 0.25
${s_{\textup{targ}}}$ Target score 90 101.25
${s_{\textup{rand}}}$ Random score 37.5
${\phi_{\textup{targ}}}$ Target bonus fee \$10
${\phi_{\textup{rand}}}$ Random bonus fee \$1
${\phi_{\textup{fix}}}$ Fixed fee \$1.5
${T_{\textup{test}}}$ Number of test rounds 5
${T_{\textup{dec}}}$ Decision inactivity timer 10s
${T_{\textup{inactive}}}$ Inactivity counter 6
Appendix B Remarks regarding departure from pre-registered analysis plan
Originally, we pre-registered the design of the experiment along with a complete plan for analysis at the Open Science Framework (Elokda and Nax, 2023). The analysis presented in the paper above constitutes a deviation from this pre-registered analysis plan. Here, for completeness, we state the hypotheses that were originally pre-registered, and discuss our motivations for deviating from the pre-registered plan.
  1. 1.
    The karma allocation is more efficient than a random allocation.
  2. 2.
    The karma allocation is more fair than a random allocation.
  3. 3.
    The efficiency of the karma allocation is within 10% of the most efficient allocation.
  4. 4.
    The karma allocation is fairer than an efficiency-maximizing but history-unaware allocation.
  5. 5.
    There is a positive correlation between the participants’ bid and urgency.
  6. 6.
    A full bidding scheme is more efficient than a binary bidding scheme.
  7. 7.
    A binary bidding scheme is more fair than a full bidding scheme.
  8. 8.
    Participants with high urgency spread achieve higher rewards than those with low urgency spread.
As regards the efficiency-related hypotheses (1), (3), (6) and (8), the results of our MWW tests strongly support hypothesis (1), weakly support hypothesis (8), and reject hypotheses (3) and (6). Note, however, that our original plan was to measure realized efficiencies using ex-post means, which was found to be overly sensitive to stochastic effects. Therefore, the analysis presented in Section 4 is based on ex-ante expected medians instead. As regards the fairness-related hypotheses (2), (4) and (7), our fairness analysis supports hypotheses (2) and (7), in particular for the high stake treatments, but not hypothesis (4). These conclusions are drawn based on comparisons of the efficiency gains across deciles, and not as measured by the negated standard deviations of final scores, as we had originally planned, which turned out to be a measure that does not produce interpretable results. Finally, the bidding behaviors portrayed in Figure 4 weakly support hypothesis (5), however, due to the bug in the logging of bids discussed in Section 4.5, we could not rigorously test this hypothesis.
Appendix C Supplementary Results
Table 4 reports the MWW test results when comparing the efficiency gains of the first versus the second halves of the experiments, per treatment. With the exception of the high stake-binary treatment, no statistically significant differences are observed.
Table 4: Mann–Whitney–Wilcoxon (MWW) tests with test statistics $U$ (and associated $p$-values in brackets) of efficiency gains in the first versus second halves of the experiments for all treatments. All treatments have $n_{\textup{karma}}=100$ samples.
Richness of scheme
  Binary   Full range Combined
Urgency process
Low stake
5 005.05005.05\,005.05 005.0
($0.9912$)
5 460.55460.55\,460.55 460.5
($0.2610$)
20 878.520878.520\,878.520 878.5
($0.4476$)
High stake
4 200.54200.54\,200.54 200.5
($0.0509$)
4 949.54949.54\,949.54 949.5
($0.9028$)
18 357.518357.518\,357.518 357.5
($0.1555$)
Combined
18 238.518238.518\,238.518 238.5
($0.1277$)
20 684.020684.020\,684.020 684.0
($0.5544$)
77 729.077729.077\,729.077 729.0
($0.4872$)
Table 5 reports the MWW test results when comparing the efficiency gains of each treatment pair as well as between combined treatments. With the exception of low stake-binary versus high stake-binary (and consequently combined low stake versus high stake), no statistically significant differences are observed.
Table 5: Inter-treatment efficiency gain MWW tests reporting test statistics $U$ (with associated $p$-values in brackets). All treatments have $n=100$ samples.
Treatment comparisons
Low stake High stake Combined
  Binary   Full range Combined   Binary   Full range Combined   Binary   Full range
Low stake
Binary   –
4 700.54700.54\,700.54 700.5
($0.4650$)
4 186.04186.04\,186.04 186.0
($0.0468$)
4 522.04522.04\,522.04 522.0
($0.2433$)
Full range
5 299.55299.55\,299.55 299.5
($0.4650$)
4 412.04412.04\,412.04 412.0
($0.1511$)
4 744.04744.04\,744.04 744.0
($0.5324$)
Combined   –   –   –   –   –
17 864.017864.017\,864.017 864.0
($0.0647$)
High stake
Binary
5 814.05814.05\,814.05 814.0
($0.0468$)
5 588.05588.05\,588.05 588.0
($0.1511$)
5 288.05288.05\,288.05 288.0
($0.4824$)
Full range
5 478.05478.05\,478.05 478.0
($0.2433$)
5 256.05256.05\,256.05 256.0
($0.5324$)
4 712.04712.04\,712.04 712.0
($0.4824$)
Combined   –   –
17 864.017864.017\,864.017 864.0
($0.0647$)
Combined
Binary   –   –   –   –   –   –   –
20 098.520098.520\,098.520 098.5
($0.9324$)
Full range   –   –   –   –   –   –
20 098.520098.520\,098.520 098.5
($0.9324$)